### v1

# Joint AP-UE Association and Power Factor Optimization for Distributed Massive MIMO

Abstract
The uplink sum-throughput of distributed massive multiple-input-multiple-output (mMIMO) networks depends majorly on Access point (AP)-User Equipment (UE) association and power control. The AP-UE association and power control both are important problems in their own right in distributed mMIMO networks to improve scalability and reduce front-haul load of the network, and to enhance the system performance by mitigating the interference and boosting the desired signals, respectively. Unlike previous studies, which focused primarily on addressing the AP-UE association or power control problems separately, this work addresses the uplink sum-throughput maximization problem in distributed mMIMO networks by solving the joint AP-UE association and power control problem, while maintaining Quality-of-Service (QoS) requirements for each UE. To improve scalability, we present an l1-penalty function that delicately balances the trade-off between spectral efficiency (SE) and front-haul signaling load. Our proposed methodology leverages fractional programming, Lagrangian dual formation, and penalty functions to provide an elegant and effective iterative solution with guaranteed convergence while meeting strict QoS criteria. Extensive numerical simulations validate the efficacy of the proposed technique for maximizing sum-throughput while considering the joint AP-UE association and power control problem, demonstrating its superiority over approaches that address these problems individually. Furthermore, the results show that the introduced penalty function can help us effectively control the maximum front-haul load for uplink distributed mMIMO systems.
Index Terms:
Distributed Massive MIMO, AP-UE Association, Power Control, Uplink Massive MIMO, Sum Spectral Efficiency
I Introduction
The increasing number of users and their demands for low latency and data-intensive services in cellular networks require adoption of technologies that can provide higher capacity, efficiency, and reliability in the networks. Massive MIMO (mMIMO) has emerged as an appealing solution in this scenario, leveraging large number of antennas to offer spatial multiplexing and channel diversity, leading to increased spectral efficiency, improved coverage, and enhanced interference management due to favorable propagation [1]. Moreover, mMIMO systems support higher user densities, making them well-suited for high-demand scenarios. However, in mMIMO based cellular systems, two key challenges need to be addressed: inter-cell interference and large variations in achievable QoS for end users [2]. Inter-cell interference stems from the overlapping coverage zones of neighboring cells, where simultaneous transmissions can cause interference to users in adjacent cells. QoS variations occur due to wide variation of signal path between UE and the Base Station (BS).
Distributed mMIMO offers a new paradigm that builds on the advantages of mMIMO in cellular networks, yet addresses the accompanying challenges [3, 4, 5]. Unlike traditional cellular networks, distributed mMIMO deploys a large number of access points (APs) throughout the coverage region. This departure from typical cell definition eliminates inter-cell interference, a persistent problem in cellular networks. Furthermore, UEs’ closeness to these distributed APs may substantially enhance the coverage probability, improving overall QoS for users.
In the original notion of distributed mMIMO [3] in the absence of traditional cell boundaries, there was a assumption that all APs simultaneously serve all UEs over the same time and frequency resources. The assumption makes the system impractical especially if the number of UEs increases dramatically. This scenario substantially increases the computational complexity and front-haul signaling load associated with Channel State Information (CSI) and data transfer, posing serious obstacles to the system’s practicality [6]. This leads to the introduction of scalable distributed mMIMO such as in [7, 6], where each UE is served by a subset of APs. This is justified as more than 95% received signal strength is concentrated in few nearby APs. Thereby, the AP-UE association is one of the important aspect of distributed mMIMO network to reduce the computation costs and front-haul load as well. Power control is another important aspect in order to achieve efficient utilization of resources in any communication network. Beyond its role in enhancing system energy efficiency, as emphasized in [5], efficient power control plays a crucial role in pro-actively minimizing interference, thereby improving the SE. Thus, the implementation of appropriate power control measures is highly desirable. So far, these power control and AP-UE association problems have been handled separately. However, jointly employing power control and AP-UE associations is more effective in mitigating interference, boosting the desired signal, and ultimately achieving superior throughput.
A significant amount of research has been undertaken on power control and AP-UE association in distributed mMIMO systems. However, it is worth noting that much of this research tends to focus entirely on either power control, as indicated by works such as [3, 5, 8, 6], or AP-UE association, as seen in [6, 9, 10]. In [5, 11], the authors focus on a downlink system, where the inherent coupling between AP-UE association and power control makes the joint analysis less complex. This intrinsic relationship provides a major advantage in the downlink scenario, facilitating joint optimization in contrast to the uplink scenario, which lacks this natural coupling between AP-UE association and power regulation. This highlights the importance of customized methodologies for extending joint optimization techniques from downlink to uplink distributed mMIMO systems. While [12] introduces a deep learning-based joint approach for power control and AP-UE association for the uplink scenario, it comes with the computational cost of extensive training and lacks the transparency associated with traditional models. In [13, 14], a novel approach was taken by formulating an optimization problem from the perspective of maximizing the minimum Signal-to-Interference-Noise ratio for joint power control and AP-UE association weights for the uplink distributed mMIMO. Notable is the inherent difficulty in allocating APs to UEs within the defined optimization challenges. This difficulty prompted the authors to suggest different methods for AP-UE scheduling.
In this work, we address the complex nature of uplink distributed mMIMO systems by formulating a mixed-integer non-convex sum-throughput maximization problem that incorporates joint AP-UE association and power control subject to the QoS requirement on each UE in terms of SE. To navigate the trade-off between Spectral Efficiency (SE) and front-haul load, we introduce an l1-penalty approach. This strategy allows for the setting up of a balanced trade-off, ensuring optimal SE while regulating the load on the front-haul of the network. Our formulation incorporates an important consideration for preserving QoS, a minimum threshold for SE for each UE thus offering equitable connectivity experience fro all the UEs. The optimization problem is solved by using the iterative approach with guaranteed convergence to a solution, by levearging the fractional programming and the Lagrangian dual formation.
Organization: The paper is organised as follows: Section II presents the system model outlining the key components and assumptions used in this study. Section II-C details the problem formulation. Section III proposes our solution and establishes its convergence. Section IV evaluates performance of the proposed scheme and benchmarks it by demonstrating its effectiveness using the numerical simulations. Finally, Section V concludes the paper by summarizing our work and discussing its future directions.
II System Model
[…]
$\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}.$ (1)
We assume that $\textbf{h}_{t,m}{\sim}\mathcal{N}_{\mathbb{C}}(0,1)$ is independent and identically distributed (i.i.d.) random vector.
II-A Uplink Pilot Training
[…]
$\textbf{y}^{pilot}_{m}=\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\mathbf{\Psi}_{t}+\textbf{n}_{m},$ (2)
[…]
$\hat{\textbf{g}}_{mt}=\frac{\beta_{mk}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t},$ (3)
Now, the mean-square of the estimate $\hat{\textbf{g}}_{mt}$ for any antenna element is given by [3]:
$\gamma_{mt}=\left(\frac{p_{p}\tau\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}p_{p}\tau\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\right)$ (4)
II-B Uplink Data Transmission
Let $x_{t}$ be the uplink payload signal transmitted from UE $t$, such that $\mathbb{E}\{|x_{t}|^{2}\}=1$. The uplink normalized signal-to-noise power is denoted by $p_{u}$. The uplink signal received at the $m^{th}$ AP is given by :
$\textbf{y}^{u}_{m}=\sqrt{p_{u}}\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\sqrt{\eta^{u}_{t}}x_{t}+\textbf{n}_{m},$ (5)
To perform detection of a signal of UE $t$ in a centralised manner, AP $m$ from the $\mathcal{M}_{t}$ send $\hat{\textbf{g}}_{mt}^{*}\textbf{y}^{u}_{m}$ to the CPU, where $\mathcal{M}_{t}\subseteq\mathcal{M}\overset{\Delta}{=}\{1,...,M\}$ is the set of APs serving the UE $t$. The received signal at CPU for the UE $t$ is [3]:
$\hat{\textbf{y}}^{u}_{t}=\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{y}^{u}_{m}]_{a},$ (6)
where [.$[.]_{a}$ represents the $a^{th}$ element of the vector [.][.][ . ]. Decomposing the $\hat{\textbf{y}}^{u}_{t}$ as in [3], results in :
$\hat{\textbf{y}}^{u}_{t}=\text{DS}_{t}\ x_{t}+\text{BU}_{t}\ x_{t}+\sum_{t^{\prime}=1,t^{\prime}\neq t}^{T}\text{UI}_{tt^{\prime}}\ x_{t^{\prime}}+n_{t},$ (7)
where
$\text{DS}_{t}\overset{\Delta}{=}\sqrt{p_{u}\eta_{t}^{u}}\mathbb{E}\left\{\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{g}_{mt}]_{a}\right\},$ (8)
$\text{BU}_{t}\overset{\Delta}{=}\sqrt{p_{u}\eta_{t}^{u}}\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{g}_{mt}]_{a}-\text{DS}_{t},$ (9)
$\text{UI}_{tt^{\prime}}\overset{\Delta}{=}\sqrt{p_{u}\eta_{t^{\prime}}^{u}}\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{g}_{mt^{\prime}}]_{a},$ (10)
$n_{t}\overset{\Delta}{=}\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{n}_{m}]_{a},$ (11)
Treating interference as effective-noise, we will get SE for the UE $t$ as :
$\text{SE}^{u}_{t}=w\text{log}_{2}\left(1+\frac{|\text{DS}_{t}|^{2}}{\mathbb{E}\{|\text{BU}_{t}|^{2}\}+\sum_{t^{\prime}=1,t^{\prime}\neq t}^{T}\mathbb{E}\{|\text{UI}_{tt^{\prime}}|^{2}\}+\mathbb{E}\{|n_{t}|^{2}\}}\right),$ (12)
where $w\overset{\Delta}{=}\left(\frac{1-\frac{L_{p}}{L_{c}}}{2}\right)$. The closed form of the SE defined in equation (12) is given as in [15]:
$\text{SE}^{u}_{t}=w\text{log}_{2}\left(1+\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right),$ (13)
where $d_{mt}$ is the indicator variable indicating the association between the UE $t$ and AP $m$, such that $d_{mt}=1$ if the AP $m$ serves the UE $t$ otherwise $d_{mt}=0$ and the $I_{t}$ is defined as :
$\text{I}_{t}\overset{\Delta}{=}\sum\limits_{t^{\prime}=1,t^{\prime}\neq t}^{T}\!\!\!\!\!A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}\left(\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}}\right)^{2}+\sum\limits_{t^{\prime}=1}^{T}\!\!\!Ap_{u}\eta_{t^{\prime}}^{u}\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!Ad_{mt}\gamma_{mt},$ (14)
II-C Problem Formulation
Our primary objective to maximize the total SE of all the UEs. Furthermore, we ensure that each UE gets minimum QoS. The minimum QoS for all the UEs is defined as the minimum SE that every UE gets. Further, to prevent the low SE providing APs from serving the UEs, we have introduced the l1-penalty in the objective function. This not only improves the energy efficiency, but also prevents the front-haul overloading. This results on the following problem formulation :
$\underset{\eta^{u}_{t},d_{mt},\textbf{D}}{\max~{}}\sum\limits_{t=1}^{T}\left(\text{SE}^{u}_{t}-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (15a)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{subject to :}\ d_{mt}\in\{0,1\},\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (15b)
$0\leq\eta^{u}_{t}\leq 1,\forall t\in\mathcal{T},$ (15c)
$\text{SE}^{u}_{t}\geq\text{SE}^{\text{QoS}}_{t},\forall t\in\mathcal{T},$ (15d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T},$ (15e)
where D is the AP-UE association matrix with each element as $d_{mt}$. $\textbf{D}_{t}$ represents the column vector of matrix of D with respect to the UE $t$. The regularization coefficient $\alpha\geq 0$ is the balancing factor between the front-haul load and SE. $\text{SE}^{\text{QoS}}_{t}$ represent the minimum QoS requirement for the UE t. The constraint (15b) represents the binary variable indicating the association between the UEs and APs. Constraint (15c) represents the uplink power control coefficient. Constraint (15d) is for the minimum QoS requirement and constraint (15e) specifies that atleast one AP must be connected to the UE. The objective function in (15a) is a non-smooth and mixed integer type non-convex function that falls into the category of the NP hard problems. Finding the optimal solution of these mixed integer type function requires high computation resources. In the upcoming sections we provide the solution of the above problems (15a) using a relaxation approach.
III Proposed Joint Solution using Relaxation Technique
The optimization problem (15a) is a mixed-integer type problem which has a very high complexity, so we convert a binary constraint (15b) into a continuous variable, thus the constraint becomes :
$d_{mt}=[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (16)
Remark 1 : The introduction of l1-penalty in the objective function serves to induce the sparsity in AP-UE association matrix D. By careful selection of the suitable trade-off factor $\alpha$, a controlled-sparsity is achieved by limiting only few APs to serve the UEs. Therefore, the matrix D has non-zero values only for the chosen AP-UE pair. After completion of AP-UE association, all entries in D undergo rounding to the nearest integer values i.e., zero and one, thus maintaining the feasibility of the optimization problem (15a). In the forthcoming result section, the SE is compared with and without rounding-off of values in D to establish the effectiveness of our proposed approach.
Then, the optimization problem (15a) can be formulate as :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}\left(1+\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (17a)
$\text{subject to :}\ d_{mt}=[0,1],\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (17b)
$0\leq\eta^{u}_{t}\leq 1,\forall t\in\mathcal{T},$ (17c)
$w\text{log}_{2}\left(1+\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)\geq\text{SE}^{\text{QoS}}_{t},\forall t\in\mathcal{T},$ (17d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T},$ (17e)
$\text{I}_{t}\geq\sum\limits_{t^{\prime}=1,t^{\prime}\neq t}^{T}\!\!\!\!\!A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})^{2}+\sum\limits_{t^{\prime}=1}^{T}\!\!\!Ap_{u}\eta_{t^{\prime}}^{u}\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\sum\limits_{m\in\mathcal{M}}\!\!\!Ad_{mt}\gamma_{mt},\forall t\in\mathcal{T},$ (17f)
Considering the non-convexity due to the ratio term inside the log in our objective function (17a), we introduce the lower-bounding auxiliary variable $\Gamma_{t}$. Then, the objective function with an additional constraint as follows :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right),$ (18a)
$\text{subject to :}\ \Gamma_{t}\leq\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}},\ t\in\mathcal{T},$ (18b)
Taking the partial Lagrangian of the objective function (18a) by considering the constraint (18b) w.r.t. $\mathbf{\Gamma}$. Then :
$L(\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma},\mathbf{\lambda})=\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}-\lambda_{t}\left(\Gamma_{t}-\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}\right)\right)$ (19)
After computing the stationary point of (19) with respect to $\mathbf{\Gamma}$, we get the optimal value of $\lambda^{*}_{t}=\frac{w^{\prime}}{1+{\Gamma^{*}_{t}}}$. where $w^{\prime}=w/log2$ and from the trivial solution optimal value of $\mathbf{\Gamma}$ is given by :
$\Gamma^{*}_{t}=\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{\text{I}_{t}}$ (20)
Substituting this $\lambda^{*}_{t}$ in (19) and using the Lagrangian dual transformation [16] then new optimization problem can be formulated as :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}-\left(\Gamma_{t}-\frac{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}\right)\right),$ (21a)
[…]
It can be observed that after substituting the optimal value of $\mathbf{\Gamma}$, as in (20), in equation (21a), the optimization problem (III) is exactly equal to solving the (III).
If we look at the optimization problem (III), we still have some non-linearities and non-convexities due to product of two variables as well as ratio term in the objective function. Firstly, we tried to handle the ratio term in the objective function using the quadratic transformation approach [16]. The quadratic transformation is performed on the objective function (21a) such that it becomes :
$\underset{\mathbf{\eta^{u}},\textbf{D},\textbf{I},\mathbf{\Gamma},\textbf{u}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)\\ -\sum\limits_{t=1}^{T}\left(\Gamma_{t}-2u_{t}\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}+u_{t}^{2}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}\right),$ (22)
The quadratic transformation $2u_{t}\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}-u_{t}^{2}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}$ is the lower bound and concave approximation of ratio term $\frac{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}}$ with respect to parameter $u_{t}$.The optimal value of u can be calculated from Equation (23). So the equation (22) is the lower bound to the equation (21a), thus making the approximation feasible.
$u^{*}_{t}\overset{\Delta}{=}\frac{\sqrt{w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+\text{I}_{t}},$ (23)
Our final optimization problem becomes:
[…]
After substituting the optimal value of u in the Equation (22), then solving optimization problem (III) is equivalent to the solving of (III).
Notice that our problem (III) is a convex optimization problem in different blocks. When fixing the D, $\mathbf{\Gamma}$ and u, then objective function of problem (24a) becomes concave w.r.t. $\mathbf{\eta^{u}}$ and I. Then, the convex-optimization problem becomes:
[…]
Similarly, upon fixing the $\mathbf{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (III) becomes concave w.r.t D and I. The convex optimization problem becomes :
[…]
So the problem can be solved following the proposed iterative approach. The detailed stepsof the proposed approach are given in Algorithm 1.
1:  Feasible Initialization: $\mathbf{\eta_{k}^{u}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall k$.
2:  Compute $I_{t}^{(0)}$ from the equation (14) respectively.
3:  repeat
4:     $i=i+1$.
5:     for all t do
6:        Compute $\Gamma_{t}$ using (20) and $u_{t}$ using (23).
7:     end for
8:     Solve optimization problem in (III) for $\mathbf{\eta_{k}^{u}}^{(i)}$ and $\textbf{I}^{(i)}$.
9:     for all t do
10:        Compute again $\Gamma_{t}$ using (20) and $u_{t}$ again using (23).
11:     end for
12:     Solve optimization problem in (III) for $\textbf{D}^{(i)}$ and $\textbf{I}^{(i)}$.
13:     Update ${\mathbf{\eta^{u}}}^{(i+1)}={\mathbf{\eta^{u}}}^{(i)}$, $\textbf{D}^{(i+1)}=\textbf{D}^{(i)}$ and $\textbf{I}^{(i+1)}=\textbf{I}^{(i)}$.
14:  until Convergence
Algorithm 1 Proposed Algorithm
The convergence proof of the proposed Algorithm 1 is given in Appendix V.
IV Performance Evaluation
We performed numerical simulations on a $1000\times 1000$ square meter geographical area with APs and user equipment UEs randomly distributed. To simulate an infinite area network, a wrapped-around approach is used [6]. Large-scale fading coefficients are modeled using a three-slope approach for path loss, incorporating an 8 dB standard deviation for shadow fading. The minimum QoS parameter is taken as 0.1 bits/s/Hz for each UE. Throughout our simulations, we keep $T=30$, $M=50\ \text{or}\ 100$, $A=4$, $\alpha=0.0005\ \text{or}\ \alpha=0.001$, $L_{c}=200$ and $L_{p}=5$, while other parameters stay exactly the same as those outlined in [3] or as explicitly stated. All these numerical results are averaged over $500$ simulations.
Figure 1: The variation of maximum fronthaul load on a AP, objective function and sum SE, when $M=50$.
Figure 2: The variation of maximum front-haul load on a AP, objective function and sum SE, when $M=100$.
Figure 1 shows the variations in maximum front-haul load [^1] The computation of front-haul load in our framework is expressed mathematically as $\sum_{t\in\mathcal{T}}d_{mt}\text{SE}^{u}_{t}\ \forall m$. Additionally, the maximum front-haul is defined as the highest front-haul load among all APs. This metric provides insight into the peak load on the front-haul infrastructure across the entire network., objective function, and sum SE for two different regularization coefficient values, with the number of antennas fixed at $50$. This highlights how the penalty function affects the network’s maximum front-haul load and total sum-throughput. Notably, the objective function value consistently remains lower than the total SE value, exhibiting a drop attributed to the penalty function. When the regularization coefficient is increased by two times, the maximum front-haul load on an AP decreases by $14\%$. This reduction is due to the increased regularization coefficient, which limits AP-UE associations. As a consequence of the lesser number of AP-UE associations, there is a reduction in sum SE by nearly $1\%$. Therefore, the proposed approach effectively decreases the maximum front-haul load while not significantly reducing overall sum throughput, and this favorable outcome corresponds to careful tuning of regularization coefficient.
Similarly, Figure 2 illustrates variability in maximum front-haul load, objective function, and sum SE for two different regularization coefficient values, with $M$ set to 100. For this scenario, when the regularization coefficient is increased two times, the maximum front-haul load and total SE show a similar trend of decrement same as in Fig. 1.
Examining Figures 1 and 2, we observe a noteworthy trend as the number of antennas ($M$) rises from 50 to 100. The maximum front-haul load increases by $14\%$ and $17\%$ with regularization coefficients ($\alpha$) of 0.0005 and 0.001, respectively. This growth can be attributed to the growing number of APs in the fixed area. As the AP count increases, more APs become closer to the each UE, resulting in a higher percentage of APs offering a strong channel or a larger large-scale fading coefficient for each UE. This increasing influence of numerous APs on the large-scale fading coefficient causes a change in the influence of the penalty function, allowing for the association of more APs with each UE. While this increased association raises the maximum front-haul load, it is important to note that increasing the number of APs results in a significant $30\%$ boost in overall sum-throughput. In essence, the system responds to the increased AP count by increasing the strength of channels across numerous APs, hence increasing the network’s overall throughput. The penalty function is a dynamic component of this process, optimizing AP-UE associations to optimize the objective function while reducing negative affects via penalties.
Figure 3: The variation of sum SE for different scenarios, when $\alpha=0.001$.
Figure 4: CDF vs per-UE SE for different scenarios, when $\alpha=0.001$.
Figures 3 and 4 illustrate the variation of sum SE and per-UE SE respectively, of the proposed solution in comparison to the following four scenarios: a) where UEs transmit at full power and are served by all APs; b) fractional power control [8]; c) where only power control is performed using the proposed algorithm; and d) where only AP-UE association is performed using the proposed algorithm.
Figure 3 shows that when AP-UE association and power control are carried out simultaneously, the achievable sum SE is the largest. The proposed solution’s sum SE surpasses that of scenarios ’a’, ’b’,’c’ and ’d’ by $16.5\%$, $4.3\%$, $1.5\%$ and $2\%$, respectively, when $M=50$. Additionally, for case with $M=100$, the proposed solution outperforms these scenarios by $11\%$, $3\%$, $1.3\%$, and $1.6\%$, respectively. This demonstrates the effectiveness of the proposed technique in optimizing the sum SE by jointly addressing the AP-UE association and power control problems. Notably, when only AP-UE association or only power control is conducted with the proposed method, the sum SE sees a improvement when compared to the other two scenarios ’a’ and ’b’. This improvement occurs while preserving the minimum QoS limitation. The proposed approach’s versatility and effectiveness are demonstrated by its ability to improve network performance even when improving just a single aspect of the system, like AP-UE association or power control.
Indeed, our proposed technique aims to enhance the sum SE. Figure 4 shows how our strategy has a significant influence on the $90\%-likely$ per-user SE. Our proposed scheme outperforms scenarios ’a’, ’b’, ’c’, and ’d’ in terms of $90\%-likely$ per-user SE by $175\%$, $22\%$, $4.7\%$, and $11.7\%$, respectively, at $M=50$. Similarly, for $M=100$, our approach outperforms the same cases by $52.3\%$, $6.7\%$, $2.1\%$, and $3.2\%$ in terms of $90\%-likely$ per-user SE. The findings illustrate an additional benefit of the method we propose, demonstrating its massive contribution to improving per-user throughput while also increasing total throughput.
Figure 5: The variation of SE with and without discretization of elements of D.
Figure 5, shows the variation of SE with and without the discretization of matrix D. The figure shows that for $m=50$, the variation of SE is negligible with and without the discretization of elements of matrix D. For the case when $m=100$, the difference in SE is mere $0.2\%$. The negligible SE variation can be attributed the fact that penalty parameter does not force elements of matrix D towards exact zero or one. This highlights the importance of penalty function in successfully forcing the value of elements of matrix D almost towards zero or one, thus our proposed solution does not require additional algorithm for AP-UE association.
Figure 6: The convergence graph of Algorithm 1, When $\alpha=0.001$, $T=30$ and $M=50$.
The convergence trajectory of Algorithm 1 is portrayed in Figure 6. The figure shows that as the number of iterations increases, so does the value of the objective function, leading up in a non-decreasing function. However, the rate of rise gradually decreases, ensuring that the rate of increase of objective function finally becomes lower than the defined tolerance of $5\times 10^{-3}$. This observation guarantees the convergence of Algorithm 1.
V Conclusion
In this work, we addresse the challenge of joint AP-UE association and power control in uplink distributed mMIMO systems by introducing an interesting l1-penalty based joint mixed-integer non-convex optimization problem. Our technique strives to maximize system throughput while adhering to minimum quality of service (QoS) requirements. For tackling this optimization problem, we reformulated the original problem and develop an iterative technique that guarantees convergence. Through numerical simulations, we demonstrate that optimizing AP-UE association and power control factor jointly results in a significant increase in the system’s sum-throughput when compared to cases where just one parameter is optimized. Another noteworthy finding is the significant decrement in the maximum front-haul load at the expense of marginal drop in the sum-throughput. Another highlight of our proposed scheme is the significant enhancement in the $90\%-likely$ per-UE SE. This enhancement demonstrates the significance and efficacy of our proposed scheme when compared to other schemes in maximizing system performance and providing a better user experience in uplink distributed mMIMO systems. Overall, our findings underline the need of optimizing AP-UE association and power regulation simultaneously, emphasizing the potential for significant throughput gains and per-user SE improvements in uplink distributed mMIMO systems.
Convergence Analysis : For the convergence analysis of the Algorithm 1, the primary goal is to prove that both optimization problem defined in (III) and (III) converges to the their respective optimal solutions. Thereafter, we prove that in each iteration of Algorithm 1, the optimization function in (24a) is non-decreasing with respect to $\mathbf{\eta^{u}}$ and D.
The feasible set of smooth concave objective function $\mathbb{F}_{1}$ (25a), articulated by the equation (25b), is closed and convex. As a result, $\mathbb{F}_{1}$ is inherently continuous and differentiable over its feasible set, implying the existence of an optimal solution for $\mathbb{F}_{1}$. The feasible set, defined by the Equation (26b), of non-smooth concave objective function $\mathbb{F}_{2}$ (26a), is closed and convex. However, $\mathbb{F}_{2}$ due to the l1-penalty lacks continuous differentiability over their feasible set. Despite this, $\mathbb{F}_{2}$ can be separated into smooth and non-smooth part. The convergence of $\mathbb{F}_{2}$ is guaranteed if gradient of the smooth part of $\mathbb{F}_{2}$ is Lipschitz continuous over its feasible set [17]. Thus the second derivative of the $\mathbb{F}_{2}$ without the non-smooth part with respect to the AP-UE association variable, $d_{mt}$ is given by :
$\frac{\partial^{2}\mathbb{F}_{1}}{\partial d_{mt}^{2}}=\sum\limits_{t=1}^{T}\left(2w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}\gamma^{2}_{mt}-\frac{P(\sum\limits_{t^{\prime}=1}^{T}A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})^{2})}{Q}-\frac{PQ^{\prime}}{Q}+\frac{PQ^{\prime 2}}{Q^{2}}\right),$ (27)
Where $P$, $Q$ and $Q^{\prime}$ are given by :
$P=w^{\prime}(1+\mathbf{\Gamma})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2},$ (28)
$Q=\sum\limits_{t^{\prime}=1}^{T}A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})^{2}+\sum\limits_{t^{\prime}=1}^{T}Ap_{u}\eta_{t^{\prime}}^{u}\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\sum\limits_{m\in\mathcal{M}}Ad_{mt}\gamma_{mt},$ (29)
$Q^{\prime}=\sum\limits_{t^{\prime}=1}^{T}A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}}+\sum\limits_{t^{\prime}=1}^{T}Ap_{u}\eta_{t^{\prime}}^{u}\gamma_{mt}\beta_{mt^{\prime}}+A\gamma_{mt}$ (30)
Notably, all the terms in the above double partial derivative (27) are bounded from above, thus the gradient of the $\mathbb{F}_{2}$ is Lipschitz continuous over its feasible set.
To prove the non-decreasing nature of the optimization function $f(\mathbf{\eta^{u}},\textbf{D})$ defined in (24a), let ${\mathbf{\eta^{u}}}^{*}$ be the optimal value of function $f$, when D is fixed. Then, for fixed D, the inequality $f({\mathbf{\eta^{u}}}^{*},{\textbf{D}}^{(i)})\geq f({\mathbf{\eta^{u}}}^{(i)},{\textbf{D}}^{(i)})$ always holds due to the concavity of the function $f$. Now, for optimal $\textbf{D}^{*}$, when $\mathbf{\eta^{u}}$ is fixed at ${\mathbf{\eta^{u}}}^{(i+1)}$, the inequality $f({\mathbf{\eta^{u}}}^{(i+1)},{\textbf{D}}^{*})\geq f({\mathbf{\eta^{u}}}^{(i+1)},{\textbf{D}}^{(i)})$ always holds as $f$ is concave for fixed D. It can be inferred that $f({\mathbf{\eta^{u}}}^{(i+1)},{\textbf{D}}^{(i+1)})\geq f({\mathbf{\eta^{u}}}^{(i)},{\textbf{D}}^{(i)})$, indicating $f$ is non-decreasing in each iteration. It makes the optimization function monotonically increasing in each iteration and also the optimization function is bounded from above, thereby guaranteeing the convergence of our proposed algorithm 1.